# Exploring Model Kinship for Merging Large Language Models

### v1

@@ 4.2 Results and Discussion
Figure 7: Weight Change.
@@
Figure 6(b) highlights the key branch of the model family tree. To investigate how merging models with low kinship helps escape local optima, we focus on the bifurcation point and analyze the weight changes: $v_{1}$ (from Model-2-1 to Model-3-1) and $v_{2}$ (from Model-2-1 to Model-3-3) in two separate experiments. The previous weight change, $v_{pre}$ (from Model-1-3 to Model-2-1), serves as a baseline. Figure 7 reveals that merging with the exploration model resulted in significant weight changes in a distinct direction, introducing novel variations into the weight space. In contrast, $v_{1}$ shows minimal weight change, as the merging effect is reduced due to the high similarity between the weights of openchat-3.5 and Model-2-1.
@@
We observe that the saturation stage of model evolution is particularly resource-intensive. In community experiments, 5 out of 14 merges in evolution path 1 resulted in an average improvement of just 0.57, while 3 out of 12 merges in evolution Path 2 yields an average improvement of 0.36. In our own experiments, applying a greedy strategy to a simple task lead to saturation after 2 out of 4 merges, with no further gains. These results indicate that human judgment and conventional stopping conditions cannot effectively halt the merging process at the optimal time. Therefore, we propose that model kinship can be used as an effective early stopping signal. When merging converges, the model kinship between top-performing models often exceeds 0.9. By halting the merging process at this point, time efficiency improves by approximately 30%, with minimal or no reduction in performance.
@@ 5 Related work
Weight averaging is one of the most widely used techniques in model merging, with its origins traced back to Utans (1996), who first applied it in neural networks to achieve performance comparable to ensemble methods. Since the 2010s, weight averaging has found numerous applications in deep neural networks, including combining checkpoints to enhance the training process (Nagarajan & Kolter, 2019; Tarvainen & Valpola, 2017; Izmailov et al., 2018; Li et al., 2023b; Stoica et al., 2023; Padmanabhan et al., 2023; Jang et al., 2023), leveraging task-specific information (Li et al., 2023a; Smith & Gashler, 2017; Ilharco et al., 2022; Izmailov et al., 2018), and parallel training of large language models (LLMs) (Li et al., 2022). Discovery of Linear Mode Connectivity (LMC) (Garipov et al., 2018; Frankle et al., 2020; Entezari et al., 2022) further expands the use of weight averaging in fusing fine-tuned models through averaging methods (Neyshabur et al., 2020; Wortsman et al., 2022). Further studies have explored optimizable weights for merging, such as Fisher-Merging (Matena & Raffel, 2022), RegMean (Jin et al., 2023), AdaMerging (Yang et al., 2024b), MaTS (Tam et al., 2024). Ilharco et al. (2023) introduce task vectors, representing the weight difference between a fine-tuned model and its base. They demonstrate that arithmetic operations on these vectors enable model editing, such as achieving multitask learning. Further research on parameter interference led to TIES (Yadav et al., 2023), which preserves important weights and reduces sign conflicts, and DARE (Yu et al., 2024), which prevents interference by randomly dropping weights. The Model Breadcrumbs (Davari & Belilovsky, 2023) show that the removal of outliers in parameters can reduce noise in model merging. Merging models with different initializations requires additional considerations. Common methods exploit the permutation symmetry of neural networks (Ainsworth et al., 2022; Tatro et al., 2020; Singh & Jaggi, 2020; Guerrero-Peña et al., 2023), using alignment techniques to mitigate the interpolation barrier (Xu et al., 2024; Navon et al., 2024). While weight averaging cannot be directly applied to models with different architectures, it can still be used to enhance feasible fusion methods. Recent work, such as FuseChat (Wan et al., 2024b), combines weight averaging with Knowledge Fusion (Wan et al., 2024a) to develop innovative fusion techniques.
@@
Recently, there have been some works exploring “model evolution”. Tellamekala et al. (2024) propose the CoLD Fusion method, showing that iterative fusion can create effective multitask models. Labonne (2024) develop a tool to automatically merge models on Hugging Face, using an “Automerge” experiment to explore metrics in the merging process. Akiba et al. (2024) introduce Evolutionary Model Merge, employing evolutionary techniques to optimize model combinations, arguing that human intuition alone cannot uncover hidden patterns in merging.
6 Conclusion and Limitations
In this paper, we introduce model kinship, the degree of similarity or relatedness between LLMs, for merging LLMs, which can help guide our selection of candidate models. We conduct comprehensive experiments to demonstrate its effectiveness in understanding the model evolution process. We further propose a new model merging strategy: Top-$k$ Greedy Merging with Model Kinship. We show that model kinship plays a crucial role in model evolution by guiding the process to escape local optima traps (in saturation stage), enabling further improvements. Additionally, we demonstrate that model kinship can detect the onset of convergence, allowing for early stopping and reducing the waste of computational resources in the merging process.
@@ 6 Conclusion and Limitations
However, there are several limitations to consider: a) The experiments in this study are conducted exclusively on models with the Mistral architecture, leaving uncertainty about the transferability of our metric and method to other architectures, such as Mamba (Gu & Dao, 2023). b) The analysis relies on open-source data from the Open Leaderboard, which is community-generated and may contain noise due to user bias. c) Correlation metrics for model kinship have not been fully explored. Other metrics may perform better than those discussed in this paper. d) The effectiveness of model kinship is demonstrated through empirical evidence. However, a theoretical framework (such as the assumptions in Appendix C) is needed to explain model evolution and model kinship more rigorously. e) Model kinship currently guides merging and improves performance limits but does not support sustained evolution. Future progress may require environmental feedback, reward models (Silver et al., 2021), as well as new architectures.
Reproducibility Statement
The experimental setup can be found in Section 4.1. All model checkpoints are available on HuggingFace, with detailed information provided in Appendices A and B.
Appendix A Details of Experiments
All merged models from these experiments are accessible through the Hugging Face Hub [^4] https://huggingface.co/datasets. The following tables cover two primary aspects:
@@ Appendix A Details of Experiments
Table 3 and 4 present detailed information on the sequential merging process. The second and third columns record the foundational models involved in each merge, while the final column lists the resulting merged models.
Table 3: Model Family tree of evolution Path 1.
@@ 14 YamShadow-7B Experiment28 YamshadowExperiment28-7B
Table 4: Model Family tree of evolution Path 2.
@@ 12 Experiment28-7B YamShadow-7B YamshadowExperiment28-7B
Table 5: Summary of Path 1 Results.
@@
Table 6: Summary of Path 2 Results.
@@
Table 7: All Sample Experiments used in distribution analysis.
@@
Appendix B Details of Model Group Selection

### v2

@@ 4.4 Discussion
Table 3: Results of merging experiments comparing the vanilla greedy strategy and our proposed approach. The first three models serve as the foundation models in both experiments. Note: The model kinship experiment was terminated at generation 5, as it has already outperformed the greedy strategy by that point.
5 Using Model Kinship to Improve Model Merging
Building on the insights from section 4, we explore how model kinship can be leveraged to improve the model merging process. Our main experiment centers on the Mistral-7B model, with detailed results presented in the main text. To further evaluate the generalization of our approach, we also conduct two supplementary experiments: one based on Llama-2 (see Appendix B) and another on a distinct task set to test robustness across different evaluation settings. Our results indicate that while greedy strategy focuses on short-term gains, it can lead to suboptimal outcomes. By integrating model kinship, we can help the strategy avoid local optima and gain further improvements.
5.1 Main Experiment Setup
For the main experiments, we follow the same settings as described in section 3, including the use of the three fine-tuned Mistral-7B variants and the evaluation on Winogrande, GSM8k, and TruthfulQA. We adopt the Top-$k$ Greedy Merging strategy as the baseline continual merging strategy.
Top $k$ Greedy Merging with Model Kinship.
We propose an enhanced merging strategy that augments the original greedy approach with an additional exploration step guided by model kinship (highlighted in blue in Algorithm 1). This approach aims to merge the best-performing model with the model that has the most distinct kinship, in order to discover potentially better solutions. In Figure 5 (b), models generated by our strategy are marked in purple, while the best-performing models are marked in yellow.
5.2 Results and Discussion
Figure 6: Weight Change between two Evolution Paths.
Figure 5 (a) illustrates the improvements in top average task performance across merging generations. Table 3 provides the model average task performance, merge gain, and model kinship for each generation, comparing the original greedy merging strategy with our kinship-based method. While both strategies achieve the multitask objective, the vanilla greedy strategy ceases to improve after Generation 2, plateauing at an average task performance of 68.72. In contrast, strategy utilizing model kinship escapes the local optima (Model-2-1) and continues to improve, reaching 69.13 by Generation 5.
@@ 5.2 Results and Discussion
Figure 5 (b) shows the main branch of the model family tree. To explore how low-kinship merges help escape local optima during saturation, we analyze weight changes: $v_{1}$ (from Model-2-1 to Model-3-1) and $v_{2}$ (from Model-2-1 to Model-3-3), with $v_{pre}$ (from Model-1-3 to Model-2-1) as baseline. Figure 6 shows that merging with the exploration model ($v_{2}$) induces significant, novel weight changes, while $v_{1}$ shows minimal change due to high similarity between openchat-3.5 and Model-2-1.
@@
Iterative merging is resource-intensive. In our main experiments, a greedy strategy saturated after 2/4 merges with no further gains. Looking back at community experiments, 5/14 merges in path 1 averaged only 0.57 improvement, and 3/12 merges in path 2 averaged 0.36. A high kinship score (PCC > 0.9) among top models may indicate convergence. Stopping merges early at high kinship generation saves  30% time with negligible performance loss.
@@ 6 Related work
Weight averaging, a widely used technique in model merging, traces its origins to Utans (1996). Since the 2010s, weight averaging has been widely applied in deep neural networks, notably for combining checkpoints to improve training stability and performance. (Nagarajan and Kolter, 2019; Tarvainen and Valpola, 2017; Izmailov et al., 2018; Li et al., 2023a; Stoica et al., 2023; Padmanabhan et al., 2023; Jang et al., 2023), leveraging task-specific information (Li et al., 2023b; Smith and Gashler, 2017; Ilharco et al., 2022; Izmailov et al., 2018), and parallel training of large language models (LLMs) (Li et al., 2022). Discovery of Linear Mode Connectivity (LMC) (Garipov et al., 2018; Frankle et al., 2020; Entezari et al., 2022) further expands the use of weight averaging in fusing fine-tuned models through averaging methods (Neyshabur et al., 2020; Wortsman et al., 2022). Further studies have explored optimizable weights for merging, such as Fisher-Merging (Matena and Raffel, 2022), RegMean (Jin et al., 2023), AdaMerging (Yang et al., 2024b), MaTS (Tam et al., 2024). Ilharco et al. (2023) introduce task vectors, representing the weight difference between a fine-tuned model and its base. Further research on parameter interference led to TIES (Yadav et al., 2023), which preserves important weights and reduces sign conflicts, and DARE (Yu et al., 2024), which prevents interference by randomly dropping weights. The Model Breadcrumbs (Davari and Belilovsky, 2023) show that the removal of outliers in parameters can reduce noise in model merging. Merging models with different initializations requires additional considerations. Common methods exploit the permutation symmetry of neural networks (Ainsworth et al., 2022; Tatro et al., 2020; Singh and Jaggi, 2020; Guerrero-Peña et al., 2023), using alignment techniques to mitigate the interpolation barrier (Xu et al., 2024; Navon et al., 2024). While weight averaging cannot be applied to models with different architectures, it can still be used to enhance feasible fusion methods. Recent work like FuseChat (Wan et al., 2024b) integrates it with Knowledge Fusion (Wan et al., 2024a) to enable novel fusion approaches.
@@
Recently, there have been some works exploring “model evolution”. Tellamekala et al. (2024) propose the CoLD Fusion method, showing that iterative fusion can create effective multitask models. Labonne (2024) develop a tool to automatically merge models on HuggingFace. Akiba et al. (2024) introduce Evolutionary Model Merge, employing evolutionary techniques to optimize model combinations.
7 Conclusion
We propose continual model merging as a framework for evolving large language models through strategic, iterative merges that yield consistent gains in generalization and task performance. To support this framework, we introduce model kinship, a metric that guides merge candidate selection and explains both performance gains and stagnation during merging. Building on this, we propose Top-$k$ Greedy Merging with Model Kinship, a strategy that uses kinship to escape local optima and achieve further improvements. Kinship also serves as an early stopping signal by detecting convergence and reducing redundant computation.
@@ 7 Conclusion
However, there are several limitations to consider: a) The experiments in this study are conducted on models with two architectures, leaving uncertainty about the transferability of our metric and method to other architectures, such as Mamba (Gu and Dao, 2023). Furthermore, scaling of tasks and candidate models requires further experimentation to understand the computational cost across various scenarios. b) The analysis is based on open source data from the Open Leaderboard, which is community-generated and may contain noise due to user bias. c) Possible correlation metrics for model kinship have not been fully explored. Other metrics may perform better than those discussed in this paper. d) The effectiveness of model kinship is demonstrated through empirical evidence. However, a theoretical framework (such as the assumptions in Appendix 15) is needed to explain model evolution and model kinship more rigorously. e) Model kinship currently guides merging and enhances performance, but does not support sustained evolution. Future progress may require environmental feedback, reward models (Silver et al., 2021), as well as new architectures.
Appendix A Details of Experiments in Main Sections
This section provides comprehensive details on the models used in the analysis of community experiments. The open merged models from these experiments are accessible through the Hugging Face Hub [^3] https://huggingface.co/datasets. The evaluation results can be accessed in the Openleaderboard [^4] https://huggingface.co/spaces/open-llm-leaderboard-old/open_llm_leaderboards The following tables cover two primary aspects:
@@ A.1 Selecting the Evolution Path
Table 4 and 5 present detailed information on the sequential merging process. The second and third columns record the foundational models involved in each merge, while the final column lists the resulting merged models.
A.2 Additional Results in Analysis
A.3 Details of Datasets Selection
In the main experiments, we utilize three task-specific benchmark datasets—Winogrande, GSM8k, and TruthfulQA—to evaluate the distinct strengths of the models. These datasets assess the following capabilities:
  • Winogrande: Evaluates the model’s commonsense reasoning.
  • GSM8k: Measures the model’s mathematical reasoning.
  • TruthfulQA: Assesses the model’s ability to identify and address human falsehoods.
@@ 14 YamShadow-7B Experiment28 YamshadowExperiment28-7B
Table 4: Model Family tree of evolution Path 1.
@@ 12 Experiment28-7B YamShadow-7B YamshadowExperiment28-7B
Table 5: Model Family tree of evolution Path 2.
@@
Table 6: Summary of Path 1 Results.
@@
Table 7: Summary of Path 2 Results.
@@
Table 8: All Sample Experiments used in distribution analysis.
Appendix B Full Evaluation Results of Main Experiments and Additional Experiments
B.1 Main Mistral-7B Experiments
Table 9 provides a detailed evaluation of the main experiments, including the results for the exploration models and their performance on specific tasks. The model kinship experiment is terminated early at generation 5, as a more promising evolution path is subsequently identified.
B.2 Additional Experiments
To assess the generalization of our strategy across different model architectures and task sets, we conduct two additional experiments.
B.2.1 Mistral-7B Experiments with a Different Task Set
We perform further evaluations using Mistral-7B, based on three distinct foundation models: MistralHermes-CodePro-7B-v1, metamath-mistral-7b, and open-chat-3.5-1210. These models are assessed on the HumanEval, GSM8k, and TruthfulQA benchmarks. The model kinship-based merging process is terminated early at generation 3, as a more promising evolution trajectory is identified.
In this task setting, model kinship-guided exploration successfully discovers models (e.g., Child3-3) that significantly outperform their respective initial baselines.
B.2.2 LLaMA-2-8B Experiments
We further evaluate our strategy on LLaMA-2-8B using three task-specific fine-tuned models. Table 11 summarizes the results of these additional experiments. The model kinship-based merging process is terminated early at generation 6 upon the discovery of a more favorable evolutionary path.
Consistent with the results from Mistral-7B, model evolution guided by model kinship continues to facilitate performance improvements beyond the capabilities of the original models.
Model TruthfulQA Winogrande GSM8K Avg. Model Kinship
MetaMath 44.89 75.77 70.51 63.72 /
Instruct 68.26 77.19 40.03 61.82 /
Open-chat 52.15 80.74 65.96 66.28 /
model-1-1-greedy 52.51 76.16 57.85 62.17 0.01
model-1-2-greedy 58.04 76.32 57.72 64.02 -0.02
model-1-3-greedy 48.96 78.69 72.86 66.84 0.05
model-2-1-greedy 50.94 80.11 75.13 68.72 0.93
model-2-2-greedy 49.78 78.93 55.72 61.47 0.57
model-2-3-greedy 52.36 78.61 52.99 61.32 0.58
model-2-exp 61.01 79.56 63.76 68.11 -0.02
model-3-1-greedy 51.95 80.51 73.31 68.59 0.95
model-3-2-greedy 49.96 79.72 73.54 67.74 0.93
model-3-3 56.95 80.25 70.00 69.06 0.24
model-3-4 54.38 78.45 73.01 68.61 0.32
model-3-exp 54.13 78.69 71.65 68.15 0.03
model-4-1-greedy 50.82 80.11 74.60 68.51 0.98
model-4-2-greedy 50.36 79.47 74.31 68.04 0.98
model-4-3-greedy 51.04 79.72 74.83 68.53 0.94
model-4-4 53.31 79.40 73.54 68.75 0.54
model-4-5 52.48 79.01 73.68 68.39 0.66
model-4-6 53.69 79.72 73.69 69.03 0.52
model-4-exp 55.16 78.53 71.80 68.49 0.48
model-5-1 54.85 79.37 73.31 69.13 0.65
model-5-2 54.78 79.40 72.86 68.98 0.65
model-5-3 53.49 79.24 73.16 68.63 0.98
model-5-exp 52.98 79.32 72.78 68.36 0.59
Table 9: Evaluation Results of Main Experiments of Mistral-7B.
Model TruthfulQA GSM8K HumanEval Avg. Model Kinship
MetaMath 44.89 70.51 17.68 44.36 /
Openchat-3.5-1210 52.15 65.96 2.44 40.18 /
MistralHermes-CodePro-7B-v1 49.67 60.88 22.56 44.37 /
child1-1-greedy 51.87 69.60 15.80 45.76 0.19
child1-2-greedy 48.04 72.78 9.15 43.32 0.08
child1-3-greedy 48.96 72.86 18.29 46.70 0.05
child2-1-greedy 50.24 71.72 12.20 44.72 0.15
child2-2-greedy 50.88 73.24 7.32 43.81 0.92
child2-3-greedy 51.15 67.32 19.51 45.99 0.34
child2-4-greedy 50.05 72.33 4.88 42.42 0.21
child2-exp 50.33 71.11 18.90 46.78 0.21
child3-1-greedy 51.47 69.22 21.34 47.34 0.73
child3-2-greedy 50.71 72.40 9.15 44.09 0.82
child3-3 49.69 74.37 21.34 48.47 0.82
child3-4 50.57 69.75 17.68 46.00 0.91
child4-1-greedy 50.56 68.46 12.20 43.74 0.79
child4-2-greedy 51.28 68.46 19.51 46.42 0.95
child5-1-greedy 51.36 68.69 20.73 46.93 0.99
child5-2-greedy 50.49 73.24 9.76 44.50 0.78
child6-1-greedy 50.50 73.24 9.15 44.30 0.78
child6-2-greedy 51.42 69.14 20.12 46.89 0.99
child7-1-greedy 51.36 68.82 20.34 46.84 0.99
child7-2-greedy 51.42 68.74 20.81 46.99 0.99
child7-3-greedy 51.44 69.15 20.44 47.01 0.99
Table 10: Evaluation Results of Additional Experiments of Mistral-7B.
Model TruthfulQA Winogrande GSM8K Avg. Model Kinship
winogrande 42.0 77.9 6.4 42.1 /
GSM8K 39.0 73.4 38.0 50.1 /
TruthfulQA 56.7 68.9 9.5 45.0 /
child1-1-greedy 40.2 79.3 34.2 51.2 0.03
child1-2-greedy 46.7 74.4 34.2 51.7 0.01
child1-3-greedy 46.1 77.1 1.9 41.7 0.01
child-2-1-greedy 44.5 78.5 36.8 53.2 0.19
child-2-2-greedy 43.7 74.0 40.4 52.7 0.45
child-2-3-greedy 38.9 77.5 37.1 51.1 0.39
child-2-exp 43.3 81.2 28.5 51.0 0.01
child-3-1-greedy 44.2 77.1 37.3 52.8 0.88
child-3-2-greedy 45.4 77.5 34.5 52.4 0.79
child-3-3-greedy 45.0 73.8 36.6 51.8 0.89
child-3-exp 45.1 78.6 30.3 51.3 0.58
child-4-1-greedy 44.4 78.5 36.8 53.2 0.95
child-4-2-greedy 44.1 75.5 40.0 53.1 0.97
child-4-exp 43.3 80.9 32.6 52.2 0.81
child-5-1-greedy 44.2 77.1 37.2 52.8 0.97
child-5-2-greedy 44.3 77.4 36.7 52.8 0.91
child-5-3-greedy 44.3 78.3 36.8 53.1 0.98
child-5-exp 44.5 78.1 32.0 51.5 0.64
child-6-1-greedy 44.5 78.5 36.8 53.2 0.99
child-6-2-greedy 44.4 78.3 36.8 53.2 0.99
child-6-3-greedy 44.3 78.3 36.8 53.1 0.99
child-6-exp 44.3 80.4 35.3 53.4 0.80
Table 11: Evaluation Results of addtional experiments of Llama-2.
Appendix C Algorithm Details for the Main Experiment
In this section, we present the detailed algorithms employed in our main experiment, along with an ablation study to validate our baseline method, Top $k$ Greedy Merging.
C.1 Algorithms
The Top-$k$ Greedy Merging algorithm aims to iteratively construct improved models through pairwise merging, guided by performance evaluation and, in the extended version, model kinship. It begins with a set of $n$ foundation models $M=\{m_{1},m_{2},\dots,m_{n}\}$. In the first step, all possible pairs of models are merged to form the first generation $G_{1}$ of merged models. These new models are added back into the candidate set $M$.
The algorithm then evaluates all models in $M$ using a task-specific evaluation function $f$ and selects the top $k$ performing models to form the working set $S$. It maintains an iterative loop that continues as long as the top-$k$ set $S$ changes between iterations, ensuring exploration continues only while performance improves. Within each iteration, all model pairs from $S$ are merged to produce the next generation of models $G_{i}$. These new models are added into $M$, and their performance is evaluated using $f$ to update $S$.
In the variant with model kinship, additional steps introduce a model exploration mechanism. This kinship-guided exploration step is designed to escape local optima by encouraging diversity in the merging path, potentially leading to models with better generalization or complementary capabilities. The algorithm terminates when the top-$k$ set $S$ stabilizes, indicating no further performance gains are observed. Each model is named according to its generation and creation order to track its origin during analysis.
C.2 Ablation Study of Greedy Strategy
The ablation study on the Greedy Strategy is conducted using the Mistral-7B architecture, following the same experimental settings outlined in the main experiments. For comparison, we employ the random-merge strategy, where models in each generation are merged with randomly selected models (excluding themselves) from the repository, as illustrated in Algorithm 2.
The following table presents the evaluation results. Each column represents:
  • Model: The name of each model. Note that the first three entries are fine-tuned foundation models used in our experiments.
  • TruthfulQA_mc2, Winogrande, GSM8K: The benchmark results for each dataset, indicating the model’s task-specific capabilities.
  • Average: The average score across all benchmarks, reflecting the model’s overall generalization performance.
  • Model Kinship: The kinship score (Here, we use cosine similarity to measure model kinship) of the parent models involved in the merge, indicating their relatedness.
  • Parent-1 and Parent-2: The names of the parent models used in the merging process.
In the random-merge strategy, the average performance in each generation fluctuates. The highest average performance achieved is 68.55, slightly lower than the 68.72 observed in the greedy experiment. While the random-merge strategy avoids convergence to local optima, it demonstrates an unstable improvement process, which can lead to unpredictable results.
Model TruthfulQA_mc2 Winogrande GSM8K Average Model Kinship
MetaMath-mistral-7B 44.89 75.77 70.51 63.72 /
Mistral-7B-Instruct-v0.2 68.26 77.19 40.03 61.82 /
Open-chat-3.5-1210 52.15 80.74 65.96 66.28 /
child1-1 52.51 76.16 57.85 62.17 0.01
child1-2 58.04 76.32 57.72 64.02 0.01
child1-3 48.96 78.69 72.86 66.84 0.03
child2-1 44.68 74.00 50.80 56.40 0.29
child2-2 49.78 78.93 55.72 61.47 0.41
child2-3 61.01 79.56 63.76 68.11 0.01
child3-1 51.52 78.23 56.71 62.15 0.84
child3-2 43.52 75.22 47.43 55.39 0.59
child3-3 54.32 78.53 72.81 68.55 0.28
child4-1 55.32 78.41 56.23 63.32 0.54
child4-2 50.53 78.42 57.65 62.20 0.86
child4-3 53.45 79.31 72.65 68.47 0.67
Table 12: Evaluation results using the random-merge strategy.
Algorithm 1 Top $k$ Greedy Merging with Model Kinship.
1:  Generate the first generation of merged models $G_{1}$ by merging each pair in set $M$, and set generation $i$ = 1.
2:  Combine the set $G_{1}$ into set $M$.
3:  Evaluate each model $m$ in set $M$, and select the top $k$ models. Denote this set as $S=\{m_{1},m_{2},\dots,m_{k}\}$.
4:  Initialize a variable $S_{\text{prev}}=\emptyset$ to store the top $k$ models from the previous iteration.
5:  while $S\neq S_{\text{prev}}$ do
6:     i++
7:     Set $S_{\text{prev}}=S$.
8:     Select each model pair from $S$. Denote this set as $P=\{p_{1},p_{2},\dots,p_{j}\}$.
9:     Merge every selected pair in set $P$ as merged model set $G_{i}=\{m_{1},m_{2},\dots,m_{j}\}$ for generation $i$, and add each merged model into set $M$.
10:     Identify the current best model $m_{best}\in S$.
11:     Identify the model $m_{f}\in S$ with the lowest model kinship to $m_{best}$ from the $G_{i-1}$ according to the similarity metric $sim(\cdot,\cdot)$.
12:     Merge $m_{f}$ with $m_{best}$ to generate a new model $m_{\text{exp}}$, and add $m_{\text{exp}}$ into set $G_{i}$ and set $M$.
13:     Evaluate each new model $m\in G_{i}$ using $f$ and update $S$.
14:     Evaluate $m_{\text{exp}}$ using $f$ and update $S$.
15:  end while
Note: The blue-highlighted steps are only executed in modified experiments incorporating model kinship-based exploration. To distinguish between different models in the subsequent experiments, each model generated in a given generation is named as model-generation-id.
Algorithm 2 Random Merge Algorithm.
0:  A set $M$ of $n$ foundation models $\{m_{1},m_{2},\dots,m_{n}\}$, Evaluation function $f$.
1:  Generate the first generation of merged models $G_{1}$ by randomly merging pairs in set $M$, and set generation $i$ = 1.
2:  Combine the set $G_{1}$ into set $M$.
3:  Evaluate each model $m$ in set $M$.
4:  Initialize a variable $S_{\text{prev}}=\emptyset$ to store the top $k$ models from the previous iteration.
5:  while $S\neq S_{\text{prev}}$ do
6:     i++
7:     Set $S_{\text{prev}}=S$.
8:     Randomly select pairs of models from $M$. Denote this set as $P=\{p_{1},p_{2},\dots,p_{j}\}$.
9:     Merge each selected pair in set $P$ to form the merged model set $G_{i}=\{m_{1},m_{2},\dots,m_{j}\}$ for generation $i$, and add each merged model into set $M$.
10:     Evaluate each new model $m\in G_{i}$ using $f$ and update $S$.
11:  end while
Appendix D Additional Analysis for Community Model Evolution
D.1 Analysis of Model Kinship Change across Merging Stages
To determine whether the discovery of increasing model kinship in model evolution paths can be generalized to the entire model evolution process, we perform an analysis of the merging stages. Given the community’s predominant use of the performance-prior strategy, we calculate model kinship among models with similar performance, simulating the selection of models at each stage. For this analysis, we randomly select 5 models from each merging stage, as delineated by boundaries identified in prior analysis - Saturation Stage ($\geq 0.75$), Improving Stage ($<$0.75 and $\geq$0.73), and Initial Merges (fine-tuned models) to form three foundation model groups, representing potential merges at different stages of model evolution.
D.2 Details of Model Group Selection